### v1

@@ Abstract
We introduce CroissantLLM, a 1.3B language model pre-trained on a set of 3T English and French tokens, to bring to the research and industrial community a high-performance, fully open-sourced bilingual model that runs swiftly on consumer-grade local hardware. To that end, we pioneer the approach of training an intrinsically bilingual model with a 1:1 English-to-French pretraining data ratio, a custom tokenizer, and bilingual finetuning datasets. We release the training dataset, notably containing a French split with manually curated, high-quality, and varied data sources. To assess performance outside of English, we craft a novel benchmark, FrenchBench, consisting of an array of classification and generation tasks, covering various orthogonal aspects of model performance in the French Language. Additionally, rooted in transparency and to foster further Large Language Model research, we release codebases, and dozens of checkpoints across various model sizes, training data distributions, and training steps, as well as fine-tuned Chat models, and strong translation models. We evaluate our model through the FMTI framework (Bommasani et al., 2023) and validate 81 % of the transparency criteria, far beyond the scores of even most open initiatives. This work enriches the NLP landscape, breaking away from previous English-centric work in order to strengthen our understanding of multilingualism in language models.
@@ 1 Introduction
Large Language Models (LLM) have taken over the Natural Language Processing (NLP) landscape in the past years. Although a few proprietary models are still considered to run ahead of the pack, open weights models such as Llama (Touvron et al., 2023a, b), Qwen (Bai et al., 2023a) or Mistral (Jiang et al., 2023, 2024) are rapidly bridging the performance gap. However, widespread industrial and research adoption of such technology remains challenging for several reasons, including the lack of transparency in the data collection and training processes, the scarcity of existing resources outside of English, and the large-scale and costly nature of existing high-performing models.
@@
Bias towards English. Although the exact training mix of most well-performing LLMs is not publicly available information, most large models are trained on very English-centric corpora. This is the consequence of the important amount of English resources compared to other languages, both in terms of data availability and benchmarks. As could be expected, all publicly available LLMs display a large performance disparity between English and non-English languages when evaluated on downstream tasks (Bandarkar et al., 2023). Moreover, cultural knowledge and biases are mechanically integrated through the training data, leading to models with greater knowledge of American events or biases (Bender et al., 2021; Santurkar et al., 2023). This puts non-English users at a disadvantage when it comes to language model usage and adoption. While the non-English NLP community has produced multilingual datasets (Soboleva et al., 2023; Laurençon et al., 2023) and models (Scao et al., 2022; Shliazhko et al., 2022) in the last few years, the available resources still largely lag behind English ones, hindering industrial adoption in international settings.
@@ 1.1 Contributions
Contribution 2: Training CroissantLLM, a truly bilingual language model. Nowadays, most models display some multilingual capabilities. For example, Bloom has been trained to be massively multilingual (Scao et al., 2022), Llama contains a minor proportion of non-English data in its training set (Touvron et al., 2023a) and Qwen included a significative portion of Chinese data (Bai et al., 2023a). [^4] Qwen-VL (Bai et al., 2023b) reports 77.3 % & 22.7 % Chinese, but no information is given for the Qwen base model. However, to our knowledge, outside of Chinese with a different alphabet(Zeng et al., 2022), no work has studied or attempted to train a multilingual model in which English is not the dominant training language.
@@
Contribution 3: FrenchBench: A novel LLM benchmark for the French Language. To evaluate models in French, we introduce a benchmark encompassing various tasks to assess factual knowledge, generative capabilities, language understanding, etc. This benchmark is constructed both from openly available datasets, as well as newly released manually annotated data. We evaluate and report results for our models as well as other models with French-speaking capabilities. [^5] Code for evaluation is currently available at https://github.com/ManuelFay/lm-evaluation-harness
Contribution 4: Releasing high-performing, inference-optimal models for the industrial community, together with a wide range of resources for the research community. The models we train are all released under open licenses. Our largest model is trained on a 3000:1 token to parameter ratio (150x longer than the Chinchilla Optimal equivalent model) leading to very strong performance in its size category. We show that model performance on downstream tasks continues to dramatically improve with lengthened pre-training runs, although model perplexity does not significantly improve. We release all checkpoints for all model sizes, as well as the exact training data seen at every step for research purposes. [^6] Training data indexing order will be released in a second stage. These models are extremely efficient to run, leading to low-latency, energy-efficient on-edge inference, even on low-resource devices such as phones or personal computers. These releases [^7] Code for dataset collection and filtering is available at https://github.com/ManuelFay/llm-data-hub. Code for model training is hosted at https://github.com/CoderPat/croissant-llm-training. Datasets and model checkpoints are available at https://huggingface.co/CroissantLLM. are motivated by a commitment to transparency to allow research and reassuring users for industrial deployment: our models comply with 81 % of criteria listed by the Foundation Model Transparency Index (Bommasani et al., 2023) (see Section 6).
@@ 2 Data
The dominance of the English language in the training data of most current models is undeniable. While multilingual models like Llama leverage some non-English data (Touvron et al., 2023a), it is only a minor part of the corpus, leading to a significant drop in performance across non-English data, with noticeable “American” bias (Santurkar et al., 2023; Navigli et al., 2023). This work aims to offset this trend by using a more balanced bilingual corpus comprising English and French, as well as additional code data. Although both languages belong to the Indo-European language family, they exhibit different morpho-syntactic structures [^8] For example, pronominal objects are placed before (resp. after) the verb in French (resp. English), both languages have different noun phrase constructions (“la commission européenne” vs. “the European commission”), etc. and French has a richer morphology. [^9] For example, English has 5 verb forms whereas French has 48, French has explicit inflections for grammatical genders, etc. However, note that only English adjectives have morphological constructions for expressing comparison (e.g. easy, easier, easiest). We refer to WALS for more details, e.g. https://wals.info/feature/21B##2/26.7/152.4 We study whether this corpus helps in reducing biases, enabling more varied knowledge sets, and unlocking non-English performance.
@@
The scrapping and processing code are available in our code base. [^11] https://github.com/ManuelFay/llm-data-hub The license information of all datasets used is given, and are permissive, including for commercial use.
@@ 2.1 French Data
Web Data. We collect internet data from various web scraps (Oscar, mC4), leveraging the CulturaX corpus (Nguyen et al., 2023) for heuristic and perplexity filtering, as well as exact and fuzzy deduplication. We also further filter the corpus to remove “binary options” scam websites, which are preponderant in the Oscar splits. In total, this represents over 363 million webpages and more than 292 billion tokens, that we split using our custom tokenizer fitted on equal amounts of French and English data. The mC4 corpus [^12] https://huggingface.co/datasets/allenai/c4 is released under the ODC-BY licence [^13] https://opendatacommons.org/licenses/by/1-0/ whereas Oscar (Abadji et al., 2022) does not claim data ownership, provides an opt-out strategy for data inclusion, and filtering metadata is released under the Creative Commons CC0 license. [^14] https://creativecommons.org/publicdomain/zero/1.0/
@@
Cultural Data. We introduce cultural data from various sources. Notably, we retrieve all Project Gutenberg (Hart, 1971) books in the French language as of October 2023, corresponding to books released in the public domain (302 million tokens). We also download and aggressively filter manuscripts and documents from the French National Library (Bibliothèque Nationale de France), and filter for documents that belong to the public domain, have undergone an OCR process, and are of high quality. [^16] Metadata is licensed under Open Etalab license https://gallica.bnf.fr/edit/und/conditions-dutilisation-des-contenus-de-gallica To filter out low-quality OCR documents, we implement custom heuristics we release within our code base. We run all documents through perplexity filters using KenLM 5-grams [^17] https://github.com/kpu/kenlm fitted on the French Wikipedia split, and discard documents with perplexity values that are too high (noisy) or too low (repetitive patterns). Thresholds are set through a manual verification process. We deliberately choose to be aggressive in our filtering to ensure only high-quality data is kept and discard the largest portion of the original corpus, keeping about 27M tokens. We choose not to keep any data from the newspaper archives, as the OCR transcription is often too noisy. More marginally, we introduce famous public domain French poems custom scrapped from a French poetry website, [^18] https://www.poesie-francaise.fr/ and run a set of podcasts through a high-quality speech-to-text model to obtain a textual transcription. This process is hard to scale and data splits from these sources are limited in quantity. Data from the OpenSubtitles [^19] https://opus.nlpl.eu/OpenSubtitles2016.php initiative is integrated, corresponding to 41.8 million tokens originating from movie subtitles. [^20] https://www.opensubtitles.org Finally, we add the French data from Wikisource collected as part of the BigScience initiative (Scao et al., 2022) and obtain 2.7 billion tokens from the process. [^21] Licensed under CC BY-SA 4.0, https://en.wikisource.org/wiki/Wikisource:Copyright_policy
Encyclopedia Data. To introduce high-quality factual data to the training corpus, we integrate the French Wikipedia split from November 2023.222222https://huggingface.co/datasets/wikimedia/wikipedia with a CC-By-SA 3.0 license This corresponds to the latest knowledge cutoff in the training data. In total, more than 2.5 million articles are used, spanning more than 2 billion tokens.
@@ 2.2 English Data
Internet Data. Similarly to the French dataset, we rely on carefully filtered content from an assortment of internet sources, including miscellaneous web pages and blogs. The filtering process includes heuristics and perplexity filtering, as well as large-scale deduplication (Soboleva et al., 2023). The SlimPajama corpus includes internet data from the CommonCrawl [^24] Common Crawl license https://commoncrawl.org/terms-of-use and C4252525ODC-BY license web scraps, as well as data sourced from Github textual content (MIT, BSD, Apache license) and the StackExchange forums. [^26] CC-By-SA 4.0 license https://archive.org/details/stackexchange
Miscellaneous. Other non-internet-based data sources are included in the SlimPajama dataset, such as scientific articles from Arxiv [^27] https://arxiv.org/, with author opt-out options and English documents from Wikipedia. [^28] CC-By-SA 3.0 license The SlimPajama dataset is also comprised of the “Books” subcorpus, obtained by downloading all book documents from Bibliotik. [^29] https://huggingface.co/datasets/the_pile_books3 Some of the documents within this last corpora have been flagged by their owner as proprietary data. We thus filter out all documents from this subcorpus to ensure no copyright claims exist within our final corpus. To compensate for this loss, we enrich the corpus with data from the open-source Project Gutenberg (Hart, 1971) English books under public domains.303030 From https://huggingface.co/datasets/pg19 with an Apache 2.0 license
@@ 2.3 Code Data
In line with most recent models (Chowdhery et al., 2022; Scao et al., 2022; Touvron et al., 2023a), we integrate code data into our training corpus. Notably, previous work shows that code data benefits natural language tasks and as such is particularly useful in data-constrained settings (Muennighoff et al., 2023). Therefore, we include 140B tokens of code data in several common programming languages. Splits and number of tokens are detailed in Table 12.
@@ 2.4 Parallel Data
Following previous work (Anil et al., 2023), we incorporate vast quantities of high-quality English-French translation pairs in order to improve the multilingual capabilities of the model (Briakou et al., 2023).
Opus. We extract subsets of sentence pairs spanning multiple domains from the OPUS corpus (Tiedemann, 2012). [^39] Free license https://opus.nlpl.eu/ Statistics are described in Table 10. In total, we include 400 million parallel sentences and about 36 billion tokens. The data is filtered through a rigorous cleaning pipeline: (1) BiFixer [^40] https://github.com/bitextor/bifixer is first used to remove duplicate data through fuzzy matching techniques; (2) BiCleaner [^41] https://github.com/bitextor/bicleaner is then used to filter data using heuristic and perplexity filters; (3) finally, the state-of-the-art NMT quality estimator CometKiwi (Rei et al., 2022) is used to keep only top quality translation pairs.
Theses. To incorporate versatile academic and scientific language, French thesis abstracts are introduced along with their author-generated English translations. This corresponds to 95000 documents and more than 80 million high-quality tokens. [^42] Data licensed under Etalab open license https://www.etalab.gouv.fr/wp-content/uploads/2017/04/ETALAB-Licence-Ouverte-v2.0.pdf
Song Lyrics. Our dataset integrates song lyrics in both French and English, scrapped from a specialized community-driven lyrics translation website. [^43] https://www.lacoccinelle.net, with opt-out options for data owners and community contributed translations under free-use license. As such, our model is trained with radically different linguistic styles (e.g. colloquialism), and the wide range of topics can help the model to capture cultural nuances. Lyrics have been translated by hand by the website community. With a total of 70k songs, we have built up a corpus of 53M tokens covering different periods (80s, 90s, 20s, etc.) and musical genres (rap, rock, jazz, etc.) To preserve colloquial expressions and cultural subtleties, we have not filtered song lyrics for explicit content. We validate the original language metadata of the songs through Google’s language-detection algorithm.

### v4

@@ Abstract
We introduce CroissantLLM, a 1.3B language model pre-trained on a set of 3T English and French tokens, to bring to the research and industrial community a high-performance, fully open-sourced bilingual model that runs swiftly on consumer-grade local hardware. To that end, we pioneer the approach of training an intrinsically bilingual model with a 1:1 English-to-French pretraining data ratio, a custom tokenizer, and bilingual finetuning datasets. We release the training dataset, notably containing a French split with manually curated, high-quality, and varied data sources. To assess performance outside of English, we craft a novel benchmark, FrenchBench, consisting of an array of classification and generation tasks, covering various orthogonal aspects of model performance in the French Language. Additionally, rooted in transparency and to foster further Large Language Model research, we release codebases, and dozens of checkpoints across various model sizes, training data distributions, and training steps, as well as fine-tuned Chat models, and strong translation models. We evaluate our model through the FMTI framework (Bommasani et al., 2023) and validate 81 % of the transparency criteria, far beyond the scores of even most open initiatives. This work enriches the NLP landscape, breaking away from previous English-centric work to strengthen our understanding of multilingualism in language models.
@@ 1 Introduction
Large Language Models (LLM) have taken over the Natural Language Processing (NLP) landscape in the past years. Although a few proprietary models are still considered to run ahead of the pack (OpenAI et al., 2023), open weights models such as Llama (Touvron et al., 2023a, b), Qwen (Bai et al., 2023a) or Mistral (Jiang et al., 2023, 2024) are rapidly bridging the performance gap. However, widespread industrial and research adoption of such technology remains challenging for several reasons, including the lack of transparency in the data collection and training processes, the scarcity of existing resources outside of English, and the large-scale and costly nature of existing high-performing models.
@@
Bias towards English. Although the exact training mix of most well-performing LLMs is not publicly available information, most large models are trained on very English-centric corpora (Touvron et al., 2023a). This is the consequence of the important amount of English resources compared to other languages, both in terms of data availability and benchmarks. As could be expected, all publicly available LLMs display a large performance disparity between English and non-English languages when evaluated on downstream tasks (Bandarkar et al., 2023). Moreover, cultural knowledge and biases are mechanically integrated through the training data, leading to models with greater knowledge of American events or biases (Bender et al., 2021; Santurkar et al., 2023). This puts non-English users at a disadvantage when it comes to language model usage and adoption. While the non-English NLP community has produced multilingual datasets (Soboleva et al., 2023; Laurençon et al., 2023) and models (Scao et al., 2022; Shliazhko et al., 2022) in the last few years, the available resources still largely lag behind English ones, hindering industrial adoption in non-English settings.
@@ 1.1 Contributions
Contribution 2: Training CroissantLLM, a truly bilingual language model. Nowadays, most models display some multilingual capabilities. For example, Bloom has been trained to be massively multilingual (Scao et al., 2022), Llama contains a minor proportion of non-English data in its training set (Touvron et al., 2023a) and Qwen included a significative portion of Chinese data (Bai et al., 2023a). [^4] Qwen-VL (Bai et al., 2023b) reports 77.3 % & 22.7 % Chinese, but no information is given for the Qwen base model. However, to our knowledge, outside of Chinese with a different alphabet (Zeng et al., 2022), no work has studied or attempted to train a multilingual model of significant scale in which English is not the dominant training language.
@@
Contribution 3: FrenchBench: A novel LLM benchmark for the French Language. To evaluate models in French, we introduce a benchmark encompassing various tasks to assess factual knowledge, generative capabilities, language understanding, etc. This benchmark is constructed both from openly available datasets, as well as newly released manually annotated data. We evaluate and report results for our models as well as other models with French-speaking capabilities. [^5] Code for evaluation is currently available at https://github.com/EleutherAI/lm-evaluation-harness [^6] Another complementary initiative has been led for French model evaluation and released concurrently in Bawden et al. (2024)
Contribution 4: Releasing high-performing, inference-optimal models for the industrial community, together with a wide range of resources for the research community. The models we train are all released under open licenses. Our largest model is trained on a 3000:1 token to parameter ratio (150 times longer than a Chinchilla Optimal 1.3B model) leading to very strong performance in its size category. We show that model performance on downstream tasks continues to dramatically improve with lengthened pre-training runs, although model perplexity does not significantly improve. We release all checkpoints for all model sizes, as well as the exact training data seen at every step for research purposes. [^7] Training data indexing order will be released in a second stage. These models are extremely efficient to run, leading to low-latency, energy-efficient on-edge inference, even on low-resource devices such as phones or personal computers. These releases [^8] Code for dataset collection and filtering is available at https://github.com/ManuelFay/llm-data-hub. Code for model training is hosted at https://github.com/CoderPat/croissant-llm-training. Datasets and model checkpoints are available at https://huggingface.co/CroissantLLM. are motivated by a commitment to transparency to allow research and reassuring users for industrial deployment: our models comply with 81 % of criteria listed by the Foundation Model Transparency Index (Bommasani et al., 2023) (see Section 6).
@@ 2 Data
The dominance of the English language in the training data of most current models is undeniable. While multilingual models like Llama leverage some non-English data (Touvron et al., 2023a), it corresponds to only a minor part of the corpus, leading to a significant drop in performance across non-English data, with noticeable “American” bias (Santurkar et al., 2023; Navigli et al., 2023). This work aims to offset this trend by using a more balanced bilingual corpus comprising English and French, as well as additional code data. Although both languages belong to the Indo-European language family, they exhibit different morpho-syntactic structures [^9] For example, pronominal objects are placed before (resp. after) the verb in French (resp. English), both languages have different noun phrase constructions (“la commission européenne” vs. “the European commission”), etc. and French has a richer morphology. [^10] For example, English has 5 verb forms whereas French has 48, French has explicit inflections for grammatical genders, etc. However, note that only English adjectives have morphological constructions for expressing comparison (e.g. easy, easier, easiest). We refer to WALS for more details, e.g. https://wals.info/feature/21B##2/26.7/152.4 We study whether this corpus helps in reducing biases, enabling more varied knowledge sets, and unlocking non-English performance.
@@
The scrapping and processing code are available in our code base. [^12] https://github.com/ManuelFay/llm-data-hub The license information of all datasets used is given, all allowing for permissive commercial use.
@@ 2.1 French Data
Web Data. We collect internet data from various web scraps (Oscar (Abadji et al., 2022), mC4 (Xue et al., 2021)), leveraging the CulturaX corpus (Nguyen et al., 2023) for heuristic and perplexity filtering, as well as exact and fuzzy deduplication. In total, this represents over 363 million webpages and more than 292 billion tokens, that we split using our custom tokenizer fitted on equal amounts of French and English data. [^13] The mC4 corpus https://huggingface.co/datasets/allenai/c4 is released under the ODC-BY licence https://opendatacommons.org/licenses/by/1-0/ whereas Oscar (Abadji et al., 2022) does not claim data ownership, provides an opt-out strategy for data inclusion, and filtering metadata is released under the Creative Commons CC0 license.https://creativecommons.org/publicdomain/zero/1.0/
@@
Cultural Data. We introduce cultural data from various sources. Notably, we retrieve all Project Gutenberg (Hart, 1971) books in the French language as of October 2023, corresponding to books released in the public domain (302 million tokens). We also download and aggressively filter manuscripts and documents from the French National Library (Bibliothèque Nationale de France), and filter for documents that belong to the public domain, have undergone an OCR process, and are of high quality. [^15] Metadata is licensed under Open Etalab license https://gallica.bnf.fr/edit/und/conditions-dutilisation-des-contenus-de-gallica To filter out low-quality OCR documents, we implement custom heuristics which we release within our code base. We run all documents through perplexity filters using KenLM 5-grams [^16] https://github.com/kpu/kenlm fitted on the French Wikipedia split, and discard documents with perplexity values that are too high (noisy) or too low (repetitive patterns). Thresholds are set through a manual verification process. We deliberately choose to be aggressive in our filtering to ensure only high-quality data is kept and discard the largest portion of the original corpus, keeping about 27M tokens. We choose not to keep any data from the newspaper archives, as the OCR transcription is often too noisy. Additionally, we introduce famous public domain French poems custom scrapped from a French poetry website, [^17] https://www.poesie-francaise.fr/ and run a set of podcasts through a high-quality speech-to-text model to obtain a textual transcription. This process is hard to scale and data splits from these sources are limited in quantity. Data from the OpenSubtitles [^18] https://opus.nlpl.eu/OpenSubtitles2016.php initiative is integrated, corresponding to 41.8 million tokens originating from movie subtitles. [^19] https://www.opensubtitles.org Finally, we add the French data from Wikisource collected as part of the BigScience initiative (Scao et al., 2022) and obtain 2.7 billion tokens from the process. [^20] Licensed under CC BY-SA 4.0, https://en.wikisource.org/wiki/Wikisource:Copyright_policy
Encyclopedia Data. To introduce high-quality factual data to the training corpus, we integrate the French Wikipedia split from November 2023.212121https://huggingface.co/datasets/wikimedia/wikipedia with a CC-By-SA 3.0 license This corresponds to the latest knowledge cutoff in the training data. In total, more than 2.5 million articles are used, spanning more than 2 billion tokens.
@@ 2.2 English Data
Internet Data. Similarly to the French dataset, we rely on carefully filtered content from an assortment of internet sources, including miscellaneous web pages and blogs. The filtering process includes heuristics and perplexity filtering, as well as large-scale deduplication (Soboleva et al., 2023). The SlimPajama corpus includes internet data from the CommonCrawl [^23] Common Crawl license https://commoncrawl.org/terms-of-use and C4242424ODC-BY license web scraps, as well as data sourced from Github textual content [^25] Only content under MIT, BSD, and Apache licenses are kept and the StackExchange forums. [^26] CC-By-SA 4.0 license https://archive.org/details/stackexchange
Miscellaneous. Other non-internet-based data sources are included in the SlimPajama dataset, such as scientific articles from Arxiv [^27] https://arxiv.org/, with author opt-out options and English documents from Wikipedia. [^28] CC-By-SA 3.0 license The SlimPajama dataset is also comprised of the “Books” subcorpus, obtained by downloading all book documents from Bibliotik. [^29] https://huggingface.co/datasets/the_pile_books3 Some of the documents within this last corpora have been flagged by their owner as proprietary data. We filter out all documents from this subcorpus, and replace them with data from the open-source Project Gutenberg (Hart, 1971) English books under public domains.303030 From https://huggingface.co/datasets/pg19 with an Apache 2.0 license
@@ 2.3 Code Data
In line with most recent models (Chowdhery et al., 2022; Scao et al., 2022; Touvron et al., 2023a), we integrate code data into our training corpus. Notably, previous work shows that code data benefits natural language tasks and can be particularly useful in data-constrained settings (Muennighoff et al., 2023). Therefore, we include 140B tokens of code data in several common programming languages. Splits and number of tokens are detailed in Table 12.
@@ 2.4 Parallel Data
Following previous work (Anil et al., 2023), we incorporate vast quantities of parallel data, in our case high-quality English-French translation pairs, in order to improve the multilingual capabilities of the model (Briakou et al., 2023).
Opus. We extract subsets of sentence pairs spanning multiple domains from the OPUS corpus (Tiedemann, 2012). [^39] Free license https://opus.nlpl.eu/ Statistics are described in Table 10. In total, we include 400 million parallel sentences and about 36 billion tokens. The data is filtered through a rigorous cleaning pipeline: (1) BiFixer (Ramírez-Sánchez et al., 2020) [^40] https://github.com/bitextor/bifixer is first used to remove duplicate data through fuzzy matching techniques; (2) BiCleaner [^41] https://github.com/bitextor/bicleaner is then used to filter data using heuristic and perplexity filters; (3) finally, the state-of-the-art NMT quality estimator CometKiwi (Rei et al., 2022b) is used to keep only top quality translation pairs.
Theses. To incorporate versatile academic and scientific language, we augment our dataset with French theses abstracts along with their author-generated English translations. This corresponds to 95000 documents and more than 80 million high-quality tokens. [^42] Data licensed under Etalab open license https://www.etalab.gouv.fr/wp-content/uploads/2017/04/ETALAB-Licence-Ouverte-v2.0.pdf
Song Lyrics. Our dataset integrates song lyrics in both French and English, scrapped from a specialized community-driven lyrics translation website. [^43] https://www.lacoccinelle.net, with opt-out options for data owners and community-contributed translations under free-use license. As such, our model is trained with radically different linguistic styles (e.g. colloquialism), and the wide range of topics can help the model to capture cultural nuances. Lyrics have been translated by hand by the website community. With a total of 70k songs, we have built up a corpus of 53M tokens covering different periods (80s, 90s, 20s, etc.) and musical genres (rap, rock, jazz, etc.) To preserve colloquial expressions and cultural subtleties, we have not filtered song lyrics for explicit content. We validate the original language metadata of the songs through Google’s language-detection algorithm.